# spacing=nonfrench

### v1

@@ Abstract.
1. Introduction
Specializing system implementations to workload characteristics and hardware can significantly improve performance and efficiency (Farshin et al., 2021; Kalia et al., 2019; Van Aken et al., 2017; Miano et al., 2022; Sriraman and Wenisch, 2018; Alipourfard et al., 2017; Karthikeyan et al., 2023; Rzadca et al., 2020; Ghasemirahni et al., 2024; Wu et al., 2024; Ghasemirahni et al., 2022; Deng et al., 2020; Sharif et al., 2018; Bhatia et al., 2004; Molnár et al., 2016; Pan et al., 2024; Sriraman et al., 2019; Dean et al., 1995; Consel and Noël, 1996). Achieving these benefits requires manual modification the system implementations and recompilation. Part of the performance benefit arises from cascading compiler optimizations, e.g. by removing a feature enabling the compiler to eliminate dead code, in turn enabling further optimizations  (Wintermeyer et al., 2020; Panchenko et al., 2019; Farshin et al., 2021). Today, system specialization for performance is manual, developer-driven, and iterative.
@@ 1. Introduction
We propose a fundamentally different approach: automated online system specialization guided by observed overall system performance. Our goal is to specialize performance critical system systems at runtime to the exact momentary workload and hardware conditions, without the human developer on the critical path. To enable this, we re-distribute tasks between ahead of time development and runtime operation. Ahead of deployment, rather than choosing a concrete set of specializations, developers specify the space of possible specializations, along with a search strategy. After deployment, our runtime iteratively generates different specialized code versions, choosing the current optimum based on observed overall system performance. This turns compile-time specialization into runtime parameter tuning, suitable for search-based auto-tuning (Van Aken et al., 2017; Sriraman and Wenisch, 2018).
To enable this, we introduce Iridescent, a framework for practical development of fast and efficient systems with online specialization. Iridescent enables both the incremental specialization of existing code-bases with minimal modifications, and in-depth design for specialization for maximizing performance. Iridescent builds on LLVM to target systems written in typical (non-managed) languages such as C(++) or Rust. Concretely, Iridescent supports the developer in separating the code-base into a specialized, performance critical core, and the remaining generic code. Iridescent then provides the developer with a specialization API to annotate possible specializations in the code thereby specifying the space of possible specializations. Additionally, Iridescent provides hooks to optionally customize JIT code generation in depth. Finally, Iridescent provides the runtime API to control the exploration of different specialized implementations as the system runs.
@@
2. Motivation & Challenges
2.1. Illustrative Example: Matrix Multiply Server
We illustrate the benefits and challenges of specialization with an example system: a server executing square matrix multiplications for clients. To optimize cache locality we use a blocked matrix multiplication (Carr and Kennedy, 1989; Wolfe, 1987, 1989). We identify one key workload parameter, $N$ the matrix size, and a key configuration parameter, $B$ the block size.
Code specialization improves performance. Both parameters offer a substantial opportunity for optimization through specialization. Unsurprisingly, we find fixing $B$ as a compile-time constant instead of leaving it a variable improves throughput by up to $6.5\times$, by enabling the compiler to unroll and vectorize the inner loops. Note that $B$ is an internal parameter that may affect performance, but any valid choice results in correct behavior with every workload. Similarly, assumptions about the workload, $N$, can also simplify the algorithm, e.g. by assuming that $N$ is a multiple of $B$ we avoid the need for copying and 0-padding partial blocks.
@@ 2.1. Illustrative Example: Matrix Multiply Server
Table 1. Optimal configurations for our block matrix multiply, across 5 hardware platforms and 3 workloads.
@@
Optimal configuration depends on workload and HW. We now compare different block sizes for different workloads (matrix sizes) on 5 different processors. As Table 1 shows, different block sizes yield optimal performance for different workload and processor combinations. Picking a fixed $B$ ahead of a practical deployment in a dynamic environment will not yield optimal performance. Moreover, even with single single-size workloads it is difficult to predict what block size will be ideal for a concrete processor.
2.2. Code Specialization is Effective but Complex
@@
There is a long line of prior work that has established the performance and efficiency benefits of specializing code, and systems code more specifically. For example, interpreted languages may generate specialized instances of functions with constant parameters and optimize them accordingly (de Assis Costa et al., 2014; Lima et al., 2020). This is a generic optimization and typically done automatically and transparently by the runtime. Other optimizations are specific to individual systems and concrete concerns, such as inlining small table lookups in software network functions as if statements (Miano et al., 2022). As most compiler optimizations, these approaches are typically guided by simple heuristics around local metrics (e.g. frequency of the same parameter value, or table size) based on developer intuition.
@@
Specialization effects often cascade. A key aspect of specialization techniques is that the effects combine. For example, PacketMill (Farshin et al., 2021) first de-virtualizes function pointers (Kohler et al., 2002) for the Click modular router (Kohler et al., 2000), and then based on this further eliminate dead code and data structure fields. The analysis for the latter optimization is impossible before de-virtualization as the virtual calls prevent the compiler from analyzing the complete packet handling code.
2.3. Specialization is Challenging in Practice
@@
(C1) Developer Complexity. A key barrier to specialization is the increased complexity for developers. Developers now need to reason about what assumptions will definitely hold and can help specialize the system for better performance. Next the developer needs to implement the required specialization and evaluate it. In practice, this frequently results in having to maintain multiple code versions. Unsurprisingly, is complex but also laborious and frequently also frustrating, since it is often an iterative hit-and-miss process.
@@
(C2) Loss of Generality. This is further complicated since specializations may hurt performance or break the system when the underlying assumptions are violated. For example, a network function may be processing 99.999% of packets with simple processing and data structures that perform vastly better with specialization. However, specializing the system to this class of packets, and thereby simplifying code and data structures breaks handling of the rare 0.001% case. The need to gracefully handle rare or unexpected cases forces developers to specialize conservatively.
@@
(C3) Optimal Specialization Choices. Finally, as we have seen above, determining the right choices to actually yield optimal overall system performance is a challenge. Modern system performance is a complex product of the emergent behaviors across all system components, it is also not feasible for developers to distill this into simple local cost models. Optimally specializing systems statically is fundamentally impossible since concrete workload and hardware parameters affect these choices, and both are dynamic in practice.
3. Approach
@@
We aim to enable practical specialization of systems for improved performance and efficiency. Our key idea to enable this given the challenges above, is to automatically explore a space of possible specializations provided by the developer at runtime based on observed overall system performance.
@@
Simple specialization with simple annotations. We first observe, that while developers struggle with determining optimal specializations and implementing a system based on them, it is much easier for developers to suggest possible specialization assumptions. We thus enable developers to annotate system code with simple annotations to this end. 2(a) shows an example for the matrix multiplication server above. Additionally, developers can also provide LLVM transformations to perform intrusive app-specific specializations, and thereby generate further candidate configurations in the overall specialization space. Crucially, we do not rely on the developer to filter or rank specializations.
@@
Guarded specialization with a fall-back. Later, when automatically instantiating specialized system versions through JIT compilation, we inject guard conditions into the code and fall back to the unspecialized version of the system code for that invocation. While triggering this guard incurs overhead, it does ensure correctness. And if the guard triggers sufficiently rarely, in combination with the benefit for the common case, the specialization may still be a net win.
@@
Online exploration guided by developer policy. Finally, we completely forgo cost models and heuristics for predicting which choice in the specialization space is optimal, and instead explore different points online as the system is running. Our guards ensure that the system always behaves correctly in this process. We rely on the developer to provide a policy for guiding this exploration. The developers control when to explore which points, manually, through existing auto-tuning solutions or by leveraging simple search strategies from our library. A key task of this policy is comparing the (application specific) overall system performance metrics, e.g. request throughput, when running different specialization points. Overall system performance metrics also implicitly factor in any overheads, e.g. triggering specialization guards for some calls. This ensures that the system converges to the optimal specialization point for the concrete combination of dynamic system, workload, and hardware conditions.
4. Iridescent Design
@@
Figure 1. Architecture of a Iridescent-specialized system.
@@ 17 Req *req = get_req(); Resp *res = new Resp();
Figure 2. MMulBlockBench microbenchmark
@@
Table 2. Iridescent API overview.
4.1. Anatomy of a Iridescent System
@@
A key component of the Iridescent mechanism is to modify code of the running system. Since Iridescent targets performance critical systems not implemented in managed languages, this is non-trivial in general. We observe though that typical performance critical systems code are almost exclusively architected with an outer loop handling a sequence of events, such as processing requests or data elements. We leverage this structure for pragmatic specialization, by requiring developers to separate the system code into two parts (Figure 1): (i) the fixed code, including initialization and outer loop; (ii) the performance-critical event handler code. The developer compiles the fixed code as before, but integrates calls to the Iridescent API (Table 2). For the handler code, the developer adds lightweight Iridescent specialization API calls and compiles to intermediate representation, LLVM IR specifically. In the fixed code, the developer obtains function pointers for the handlers through the Iridescent policy API, and calls these as before. During dynamic code generation, Iridescent will link handlers against symbols in the fixed code, to enable calls and accesses to global state. Since Iridescent will repeately rebuild handlers, Iridescent requires that the handler code does not include definitions of global variables that must be preserved. The developer has to move all global state to the fixed code, and reference from the handlers. Figure 7 illustrates this with our matrix multiplication server running example.
4.2. Defining Specialization Space
@@
The developer defines the specialization space through specialization point annotations in the handler code. At runtime, Iridescent replaces these annotations with specialized code, instrumentation, or disables them, guided by the policy. Iridescent offers three types of specialization points:
@@
Value Specialization Point. A value specialization point signals to Iridescent that the handler code could be specialized to a constant for the wrapped expression. spec_enum indicates the value could be one of the specified values; spec_range instead specifies a range; while spec_generic leaves it to the policy to determine possible values.
@@
Assumption Specialization Point. This point provides a possible specialization assumption to Iridescent. The condition is a boolean predicate that Iridescent can provide to compiler optimizations to further optimize, through LLVM’s builtin llvm.assume. Crucially, the assumption need not always hold, unlike for the LLVM intrinsic where incorrect behavior results, Iridescent catches this with its specialization guards.
Custom User-Defined Specialization Point. These specialization points, annotated through spec_custom_*, invoke developer-defined specializations registered in the fixed code through add_custom_spec. An example is generating an if-else chain as a special-case fast-path for an expensive data structure lookup or computation (§ 5.1).
4.3. Defining Specialization Policy
The developer implements the specialization policy in the system fixed code using the Iridescent policy API. Through the policy, the fixed code decides what to specialize and when. Iridescent provides a simple periodic exhaustive search strategy for simple cases as a library routine, with a simple call-back for providing the strategy with the system performance metric. But we also found in our evaluation there are typically system-specific insights to more intelligently decide when to trigger exploration, and how to prioritize different configurations. As a result, we expect that most systems will implement custom strategies maximizing overall system performance, using the Iridescent building blocks.
2(b) shows how through a simple specialization policy, developers can control the exploration and selection of specializations. In line 4, the spec_space API returns the specialization space generated by the annotations in the handler code. Here, we then combine this space, with other specializations through a cartesian product to obtain a complete set of specializations. In lines 4-8, the fixed code automatically tries out the different combinations in the set of specializations and chooses the best performing combination. To choose the best, the specialization policy uses the target end-to-end performance metric, in this case the throughput. Here, we then trigger re-exploration at a fixed-time interval to adapt to workload changes.
The specialization policy provides developers fine-grained control for specialization. This also includes enabling instrumentation for specialization points to dynamically identify opportunities for specialization. For example, the policy for the matrix multiplication server may specify that the matrix size, $N$, should only be specialized if $\geq$70% of the workload has the same value. The developer can also optionally configure custom optimization passes through the customize_opts API to further modify code generation.
4.4. Specialization Runtime
The specialization runtime has two key components: (i) the specializer generating the specialized code by applying selected specializations; (ii) the JIT compiling and optimizing the modified specialized code and making it available to the fixed code.
4.4.1. Specializer
The specializer generates the specialized intermediate representations for the handler code. It operates at function granularity, and generates versions of functions with specialization points replaced depending on the chosen specialization configuration.
For each specialization point, Iridescent performs the following actions, depending on the configuration the policy supplied for this point. If the policy marks a point as disabled, the Iridescent specializer simply removes the annotation and replaces it by the original value, or skips over assumption points. To specialize an enabled value specialization point, Iridescent replaces the specialization point annotation in the handler code with the constant value supplied by the policy. By default, the specializer will also insert a specialization guard, which the developers may explicitly disable. To specialize a custom specialization point, the specializer replaces the specialization point with the custom source code provided by the previously registered handler for that custom specialization point type. The recompilation process then enables typical compiler optimization passes to take effect that now operate with the new constants.
Instrumentation. Some specialization policies benefit from collecting runtime data to generate possible specialization configurations. To support such data collection, the specializer can optionally enable instrumentation for each specialization point. For a specialization point with instrumentation enabled, Iridescent additionally generates code for collecting and storing the observed actual values. The policy retrieves this information included in the result of the spec_space call.
4.4.2. JIT
Compiling the specialized code. The specialization runtime adds the specialized code generated by the specializer to the JIT. The JIT compiles the code and runs the developer-specified pipeline of optimization passes, including any custom developer-provided optimization passes, to generate an optimized version of the specialized code. By default, the JIT applies the default O3 optimization pipeline if the developer has not provided a specific pipeline. In addition, the JIT also keeps a copy of the original, non-specialized version of the specialized function. This generic function acts as the fallback option for situations where the specialized function is not applicable.
Using the specialized code. To use the specialized handlers, the fixed code obtains function pointers from the specialization runtime. The fixed code then uses to invoke the specialized handlers. If no specialization has been enabled the handlers default to the non-customized generic version of the functions. For simplicity, the fixed code only needs to do this once at the very start of execution. The JIT creates a trampoline function which calls the most recent specialized version of the function. The trampoline function is stored at a fixed address and does not change across runtime updates.
4.4.3. Correctness with Specialization Guards
Iridescent by default inserts a runtime check called a specialization guard for an enabled specialization point. The guard can be disabled by the policy. The specialization check ensures that the condition for using the specialized version of the code holds during execution. For this, specializer inserts code to perform an early exit from the specialized version of the code on check failure by throwing an exception. The JIT’s trampoline function catches the thrown exceptions and re-routes the control flow to the original non-specialized version of the function transparently without exposing the exception to the fixed code.
Restoring state and side effects. To ensure correct restoration of the state, the specializer will call a user-defined cleanup function for reversing any side-effects before throwing the exception. It is critical to note here that not all side-effects are reversible (e.g. sending a packet to a neighbor), so Iridescent performs a best-effort clean-up.
4.5. Prototype Implementation
We have implemented Iridescent in $5K$ lines of C++. Our implementation uses the LLVM IR and JIT (Lattner and Adve, 2004) to generate the specialized code at runtime. The specializer is implemented as a set of LLVM transformation passes that operate on LLVM IR of the instrumented handler code.
5. Case Studies
For our evaluation we use four open-source systems and libraries as our evaluation target and the MMulBlockBench microbenchmark. We explain each system along with the changes we made for integrating Iridescent below.
@@ 5. Case Studies
In addition to the benchmarks, we also implemented a generic version of the hot key specialization of Morpheus (Miano et al., 2022) called the fast-path specialization. The fast-path specialization works in two phases: (i) the instrumentation phase: Iridescent inserts instrumentation code that samples (sampling rate selected by the user) invocations of the target function to find the most popular inputs to the function along with their calculated outputs; (ii) the specialization phase: Iridescent updates the target function to have a specialized if-else chain where the top-N inputs along with their outputs are converted into a series of if-else checks to avoid paying the cost of re-computation for the heavy-hitting inputs. The value of N maybe user provided or it could also be configured as a runtime constant specialization point. If the input does not match any of the branches in the if-else chain, then the computation simply defaults to the generic version.
5.1. LibLPM

### v2

@@ Abstract
1 Introduction
Specializing system implementations to workload characteristics and hardware can significantly improve performance and efficiency [27, 39, 82, 53, 78, 4, 41, 70, 31, 86, 30, 24, 72, 8, 55, 60, 77, 22, 18, 16, 76]. Achieving these benefits requires manual modification of the system implementations and recompilation. Part of the performance benefit arises from cascading compiler optimizations, e.g. by removing a feature enabling the compiler to eliminate dead code, in turn enabling further optimizations  [83, 61, 27]. Today, system specialization for performance is manual, developer-driven, and iterative.
@@ 1 Introduction
We propose a fundamentally different approach: automated online system specialization guided by observed overall system performance. Our goal is to specialize performance critical system systems at runtime to the exact momentary workload and hardware conditions, without the human developer on the critical path. To enable this, we re-distribute tasks between ahead of time development and runtime operation. Ahead of deployment, rather than choosing a concrete set of specializations, developers specify the space of possible specializations, along with a search strategy. After deployment, our runtime iteratively generates and measures the performance of different specialized code versions, choosing the currently optimal specialization based on observed overall system performance. This allows compile-time specializations at runtime, suitable for search-based auto-tuning [82, 78, 75].
To enable this, we introduce Iridescent, a framework for practical development of fast and efficient systems with online specialization. Iridescent enables both the incremental specialization of existing code-bases with minimal modifications, and in-depth design for specialization for maximizing performance. Iridescent builds on LLVM and targets systems written in typical (non-managed) languages such as C(++) or Rust that are architected around an event-loop. Concretely, Iridescent requires developers to separate the code-base into a specialized, performance critical handler core, and the remaining outer event loop code. Iridescent then provides the developer with a specialization API to annotate possible specializations in the code thereby specifying the space of possible specializations. Additionally, Iridescent provides hooks to optionally customize JIT code generation in depth. Finally, Iridescent provides a runtime API to control the exploration of different specialized implementations as the system runs.
@@
2 Motivation & Challenges
We illustrate the benefits and challenges of specialization with an example system: a server executing square matrix multiplications for clients. To optimize cache locality we use a blocked matrix multiplication algorithm [11, 84, 85]. We identify one key workload parameter, $N$, the matrix size, and a key configuration parameter, $B$ the block size.
Code specialization improves performance.
Both parameters offer a substantial opportunity for optimization through specialization. Unsurprisingly, we find fixing $B$ as a compile-time constant instead of leaving it a variable improves throughput by up to $6.5\times$, by enabling the compiler to unroll and vectorize the inner loops. Note that $B$ is an internal parameter that may affect performance, but any valid choice results in correct behavior with every workload. Similarly, assumptions about the workload, $N$, can also simplify the algorithm, e.g. by assuming that $N$ is a multiple of $B$ we avoid the need for copying and 0-padding partial blocks.
@@ 2 Motivation & Challenges
Table 1: Optimal value of block size, $B$, for our block matrix multiply, across 5 hardware platforms and 3 workloads.
Optimal configuration depends on workload and HW.
@@
We now compare different block sizes for different workloads (matrix sizes) on 5 different processors. As Table 1 shows, different block sizes yield optimal performance for different workload and processor combinations. Picking a fixed $B$ ahead of a practical deployment in a dynamic environment will not yield optimal performance. Moreover, even with single single-size workloads it is difficult to predict what block size will be ideal for a concrete processor.
2.1 Specialization is Effective but Complex
@@
There is a long line of prior work that has established the performance and efficiency benefits of specializing code, and systems code more specifically. For example, interpreted languages may generate specialized instances of functions with constant parameters and optimize them accordingly [21, 49]. This is a generic optimization and typically done automatically and transparently by the runtime. Other optimizations are specific to individual systems and concrete concerns, such as inlining small table lookups in software network functions as if statements [53]. Like most compiler optimizations, these approaches are typically guided by simple heuristics around local metrics (e.g. frequency of the same parameter value, or table size) based on developer intuition.
Specialization effects often cascade.
@@
A key aspect of specialization techniques is that the effects combine. For example, PacketMill [27] first de-virtualizes function pointers [43] for the Click modular router [44], and then based on this further eliminates dead code and data structure fields. The analysis for the latter optimization is impossible before de-virtualization as the virtual calls prevent the compiler from analyzing the complete packet handling code.
2.2 Specialization is Challenging in Practice
(C1) Developer Complexity.
@@ 2.2 Specialization is Challenging in Practice
A key barrier to specialization is the increased complexity for developers. Developers now need to reason about what assumptions will definitely hold and can help specialize the system for better performance. Next the developer needs to implement the required specialization and evaluate it. In practice, this frequently results in having to maintain multiple code versions. Unsurprisingly, this is not only complex but also laborious and frustrating, since it is often an iterative hit-and-miss process.
(C2) Loss of Generality.
@@
Specializations may hurt performance or break the system when the underlying assumptions are violated. For example, a network function may process 99.999% of packets if the processing loop and data structures are specialized in a specific way. However, specializing the system to this class of packets, and thereby simplifying code and data structures may break handling of the rare 0.001% cases. The need to gracefully handle rare or unexpected cases forces developers to specialize conservatively.
(C3) Optimal Specialization Choices.
@@
Modern system performance is a complex product of the emergent behaviors across all system components, it is also not feasible for developers to distill this into simple local cost models. Optimally specializing systems statically is fundamentally impossible since concrete workload and hardware parameters affect these choices, and both are dynamic in practice.
3 Approach
Specialization Effect
Value Convert runtime variables to constants; triggers const. prop., unrolling, vectorization, ….
Assumption Add boolean assumptions; triggers dead code elimination, branch elimination, …
Fast-Path Add if-else based early exits with previously computed/cached results.
Custom Code Generate custom code at runtime. e.g. dynamic LPM matching via nested branches.
Table 2: Iridescent specialization types and their effects.
@@ 3 Approach
We enable practical specialization of systems for improved performance and efficiency. Our key idea to enable this given the challenges above, is to automatically explore a space of possible specializations provided by the developer at runtime based on observed overall system performance.
Simple specialization with simple annotations.
@@
We first observe that, while developers struggle with determining optimal specializations and implementing a system based on them, it is much easier for developers to suggest possible specialization assumptions. We thus enable developers to annotate system code with simple annotations to this end. 2(a) shows an example of the annotations for the matrix multiplication server above and how developers can configure specializations at runtime. Additionally, developers can also provide LLVM transformations to perform intrusive app-specific specializations, and thereby generate further candidate configurations in the overall specialization space. Crucially, we do not rely on the developer to filter or rank specializations. Table 2 shows the broad class of specializations supported by Iridescent and their corresponding effects.
Guarded specialization with a fall-back.
@@
Later, when automatically instantiating specialized system versions through JIT compilation, we inject guard conditions into the code to use the specialized version of the code when applicable. If the specialized version is not applicable, the guard conditions ensure that we fall back to the unspecialized version of the system code for that invocation. While triggering this guard incurs overhead, it does ensure correctness. If the guard triggers sufficiently rarely, in combination with the benefit for the common case, the specialization is still a net win.
Online exploration guided by developer policy.
@@
Finally, we completely forgo cost models and heuristics for predicting which choice in the specialization space is optimal, and instead explore different points online as the system runs. Our guards ensure that the system always behaves correctly in this process. We rely on the developer to provide a policy for guiding this exploration. The developers control when to explore which points, manually, through existing auto-tuning solutions or by leveraging simple search strategies from our library. A key task of this policy is comparing the (application specific) overall system performance metrics, e.g. request throughput, when running different specialization points. Overall system performance metrics also implicitly factor in any overheads, e.g. triggering specialization guards for some calls. This ensures that the system converges to the optimal specialization point for the concrete combination of dynamic system, workload, and hardware conditions.
4 Iridescent Design
@@
Figure 1: Architecture of a Iridescent-specialized system.
@@ 17 Req *req = get_req(); Resp *res = new Resp();
Figure 2: MMulBlockBench microbenchmark
@@
Table 3: Iridescent API overview.
4.1 Anatomy of a Iridescent System
System Requirements for Iridescent.
Iridescent targets performance critical systems not implemented in managed languages. We observe that typical performance critical systems code is almost exclusively architected with an event loop structure – an outer loop handling a sequence of events, such as processing requests or data elements, and an inner handler that processes the event. While Iridescent is designed to specialize these systems in varying dynamic environments, for these systems to fully gain the benefits of Iridescent, these systems should have stability in the workload and operating conditions for a sufficient period of time. Conditions should at least remain stable for the time Iridescent takes to re-compile a new specialized variant (typically 10–400 ms, see § 6.4).
Isolating performance critical code for specialization.
@@ 4.1 Anatomy of a Iridescent System
Since Iridescent targets performance critical systems not implemented in managed languages, modifying code of the running system is non-trivial in general. We leverage the event-loop structure for pragmatic specialization, by requiring developers to separate the system code into two parts (Figure 1): (i) the fixed code, including initialization and the outer loop; (ii) the performance-critical event handler code. The developer compiles the fixed code as before, but integrates calls to the Iridescent API (Table 3). In the handler code, the developer adds lightweight Iridescent specialization API calls and compiles to LLVM intermediate representation. In the fixed code, the developer obtains function pointers for the handlers through the Iridescent policy API, and calls these as before. Figure 2 illustrates this with our matrix multiplication server running example.
Handling Global State.
During dynamic code generation, Iridescent will link handlers against symbols in the fixed code, to enable calls and accesses to global state. Since Iridescent repeatedly rebuilds handlers, Iridescent requires that the handler code does not include definitions of global variables that must be preserved. The developer has to move all global state to the fixed code, and reference it from the handlers.
4.2 Defining Specialization Space
@@
The developer defines the specialization space through specialization point annotations in the handler code. At runtime, Iridescent replaces these annotations with specialized code, instrumentation, or disables them, guided by the policy. Iridescent offers four types of specialization points:
Value Specialization Point.
@@
A value specialization point signals to Iridescent that the handler code could be specialized to a constant for the wrapped expression. spec_enum indicates the value could be one of the specified values; spec_range instead specifies a range of integers; spec_generic leaves it to the policy to determine possible values.
Assumption Specialization Point.
@@
This provides a possible specialization assumption to Iridescent. The condition is a boolean predicate that Iridescent can provide to compiler optimizations to further optimize, through LLVM’s builtin llvm.assume. Crucially, the assumption need not always hold. Unlike for the LLVM intrinsic where incorrect behavior results, Iridescent catches this with its specialization guards.
Fast-Path Specialization Point.
@@
This is a generic version of the hot key specialization of Morpheus [53]. It works in two phases: (i) the instrumentation phase: Iridescent inserts instrumentation code that samples (sampling rate selected by the developer) invocations of the specialized function to find the most popular inputs to the function along with their calculated outputs; (ii) the specialization phase: Iridescent updates the specialized function to have a specialized if-else chain where the top-N inputs along with their outputs are converted into a series of if-else checks to avoid paying the cost of re-computation for the heavy-hitting inputs. The value of N may be developer-provided or could be configured as a runtime constant that can be further specialized. If the input does not match any of the branches in the if-else chain, the computation simply defaults to the generic version.
Custom User-Defined Specialization Point.
These specialization points, annotated through spec_custom_*, invoke developer-defined specializations registered in the fixed code through add_custom_spec. An example is generating an if-else chain as a special-case fast-path for an expensive data structure lookup or computation (§ 5.2).
4.3 Defining Specialization Policies
The developer implements the specialization policy in the system’s fixed code using the Iridescent policy API. The policy decides what, when, and how Iridescent should specialize.
2(b) illustrates how developers can control the exploration and selection of specializations. In line 4, the spec_space API returns the specialization space generated by the annotations in the handler code. We combine this space with other specializations through a Cartesian product to obtain a complete set of specializations. In lines 4–8, the fixed code automatically tries out the different combinations in the set of specializations and chooses the best performing combination. To choose the best, the specialization policy uses the target end-to-end performance metric, in this case the throughput. We trigger re-exploration at a fixed-time interval to adapt to workload changes.
The developer can also optionally configure custom optimization passes through the customize_opts API to interpose on and customize code generation. Developers may also choose to enable instrumentation for select specialization points to dynamically identify opportunities for specialization. For example, the policy for the matrix multiplication server could detect the most frequent matrix size $N$, and only decide to specialize if $\geq$70% of requests have this value.
Exploring the specialization space.
Iridescent provides a periodic exhaustive search strategy for simple cases as a library routine, with a simple call-back for providing the strategy with the system performance metric. For more complex cases, Iridescent integrates with OPPerTune [75] to leverage its state-of-the-art search strategy for configuration tuning of web services. Additionally, Iridescent also provides a thin interface for developing custom search strategies. We expect most systems will either utilize existing auto-tuning or implement custom strategies maximizing overall system performance, using the Iridescent building blocks.
4.4 Specialization Runtime
The specialization runtime has two key components: (i) the specializer generating code by applying selected specializations; (ii) the JIT compiling and optimizing the modified specialized code and making it available to the fixed code.
4.4.1 Specializer
The specializer generates the specialized version of the handler code. It operates at function granularity, and generates versions of functions with specialization points replaced depending on the chosen specialization configuration.
For each specialization point, Iridescent performs the following actions, depending on the configuration the policy supplied for this point. If the policy marks a point as disabled, the Iridescent specializer simply removes the annotation and replaces it with the original value, or skips over the point if it is an assumption point. To specialize an enabled value specialization point, Iridescent replaces the specialization point annotation in the handler code with the constant value supplied by the policy. By default, the specializer will also insert a specialization guard, which the developers may explicitly disable. For a custom specialization point, the specializer replaces the specialization point with the custom source code provided by the previously registered handler for that point type. The recompilation process then enables standard compiler optimization to take advantage of the newly inserted constants or code.
Instrumentation.
Some specialization policies benefit from collecting runtime data to generate possible specialization configurations. To support such data collection, the specializer can optionally enable instrumentation for each specialization point. For a specialization point with instrumentation enabled, Iridescent additionally generates code for collecting and storing the observed actual values. The policy retrieves this information included in the result of the spec_space call.
4.4.2 JIT
Compiling the specialized code.
The specialization runtime adds the specialized code generated by the specializer to the JIT. The JIT compiles the code and runs the developer-specified pipeline of optimization passes, including any custom developer-provided optimization passes, to generate an optimized version of the specialized code. By default, the JIT applies the default O3 optimization pipeline if the developer has not provided a specific pipeline. In addition, the JIT also keeps a copy of the original, non-specialized version of the specialized code. This generic code serves as fallback for situations where the specialized code is not applicable.
Using the specialized code
. To use the specialized handlers, the fixed code obtains function pointers to specialized code from the specialization runtime. The fixed code then uses these pointers to invoke the specialized handlers. If no specialization has been enabled the handlers default to the non-specialized versions of the functions. For simplicity, the fixed code only does this once at start of execution. The JIT creates a trampoline function which calls the most recent specialized version of the function. The trampoline function is stored at a fixed address and does not change across specialization iterations.
4.4.3 Correctness through Specialization Guards
Iridescent inserts a runtime check called a specialization guard for each enabled specialization point. The policy can disable this guard for individual points. The specialization guard ensures that the specialization conditions for generated code holds during execution. For this, the specializer inserts an early exit from the specialized version of the code on guard failure by throwing an exception. The Iridescent trampoline function catches these exceptions and re-routes control to the original non-specialized version of the function transparently without exposing the exception to the fixed code.
Restoring state and side effects.
To ensure correct restoration of the state, the specializer will call a developer-defined cleanup function for reversing any side-effects before throwing the exception. It is critical to note here that not all side-effects are reversible (e.g. sending a packet to a neighbor), so Iridescent leaves the developer with the necessary control.
4.5 Prototype Implementation
We have implemented Iridescent in $5.5K$ lines of C++. Our implementation uses the LLVM IR and JIT [45] to generate the specialized code at runtime. The specializer is implemented as a set of LLVM transformation passes that operate on the LLVM IR of the instrumented handler code.
5 Case Studies
We evaluate Iridescent with four open-source systems and libraries and the MMulBlockBench microbenchmark. We explain each system along with the changes we made for integrating Iridescent below.
5.1 Tiling Benchmarks
MMulBlockBench.
We implement a blocked/tiled version of matrix multiplication algorithm [11, 84, 85]. To integrate this benchmark with Iridescent, we convert the tile size, $B$, as a value specialization point.
CloverLeaf.
We use a C-version of the CloverLeaf application from the SPEC ACCEL2023 benchmark [19]. In this application, we implement tiled-versions of the various mathematical kernels, and use value specialization points for the tile sizes, $T_{i}$ and $t_{J}$. In total, this application has a 24 unique specialization points, each of which that can take 7 unique valeus resulting in a total of $7^{24}$ possible configurations.
5.2 LibLPM